.The concentration of a solution of EDTA was determined by standardizing against a solution of Ca2+ prepared from the primary standard CaCO3. A 0.4071-g sample of CaCO3 was transferred to a 500-mL volumetric flask, dissolved using a minimum of 6 M HCl, and diluted to volume. A 50.00-mL portion of this solution was transferred into a 250-mL Erlenmeyer flask and the pH adjusted by adding 5 mL of a pH 10 NH3–NH4Cl buffer containing a small amount of Mg2+–EDTA. After adding EBT as a visual indicator, the solution was titrated with the EDTA, requiring 42.63 mL to reach the end point. Report the molar concentration of the titrant


1
Expert's answer
2021-08-26T01:08:33-0400

CaCO3(s) + 2HCl(aq) → Ca2+(aq)+2Cl1(aq)+H2O(l) + CO2(g)


Mole of Ca2+ in CaCO3 = 0.4071g/100g/mol

= 0.00407mol

Molarity of Ca²+ = 0.00407mol/0.5L

= 0.00814M


From Dilution theory

C1 V1 = C2V2

That's conc of the solution in 50ml part

0.00814M × 0.5L= C1 × 0.05L

C1 = 0.0814M


Mole of Titrant= 0.0814M/50ml

= 1.628mol


Molar concentration of the Titrant= 1.628mol/42.63 mL

= 0.0381M

QED